delivery service in Metro Manila has a mean delivery time of 55 minutes with a standard deviation of 20 minutes. Assuring that the distribution is normally distributedthen the probability that the food will be delivered for less than 25 minutes is Blank 1. Express your answer in percentage: rounded-off to two 2) decimal places.
1
Expert's answer
2020-10-21T15:38:57-0400

"P(x<25)=P(Z<\frac{25-55}{20})"

"=P(Z<-1.5)=0.0668"

=6.68%